A gas has a volume of 4.24 L at 0 ∘C . What final temperature, in degrees Celsius, is needed to change the volume of the gas to each of the following, if n and P do not change?

1. 1.64L express your answers to three significant figures in degrees celsius

2. 1260mL express your answers to three significant figures in degrees celsius

3. 10.1L express your answers to three significant figures in degrees celsius

4. 59.9mL express your answers to three significant figures in degrees celsiu
Answer #1

1.

Initial                                        Fianl

V1 = 4.24L                              V2 =1.64L

T1 = 0+273 = 273K               T2 =

       V1/T1   = V2/T2

         T2      = V2T1/V1

                     = 1.64*273/4.24   = 105.6K

0C= 105.6-273    = -1670C

2.

Initial                                        Fianl

V1 = 4.24L                              V2 =1260ml   = 1.260L

T1 = 0+273 = 273K               T2 =

       V1/T1   = V2/T2

         T2      = V2T1/V1

                   = 1.260*273/4.24    = 81K

0C= 81-273    = -1920C

3.

Initial                                        Fianl

V1 = 4.24L                              V2 =10.1L

T1 = 0+273 = 273K               T2 =

       V1/T1   = V2/T2

         T2      = V2T1/V1

                   = 10.1*273/4.24    = 650K

0C= 650-273    = 3770C

4.

Initial                                        Fianl

V1 = 4.24L                              V2 =59.9ml   = 0.0599L

T1 = 0+273 = 273K               T2 =

       V1/T1   = V2/T2

         T2      = V2T1/V1

                   = 0.0599*273/4.24    = 3.86K

0C= 3.86-273    = -2690C